OpenStudy (anonymous):

Divide and simplify. Assume that all variables represent nonnegative numbers: √(16/36) =

OpenStudy (anonymous):

divide both numerator and denominator by 4 and see what you get...then proceed further

OpenStudy (anonymous):

4 and 9

OpenStudy (anonymous):

the common denominator is 4

OpenStudy (anonymous):

now use the rule \[\sqrt{\frac{ m }{ n }}=\frac{ \sqrt{m} }{ \sqrt{n} }\]

OpenStudy (anonymous):

friends pls do not give out the answer...

OpenStudy (anonymous):

What I find helpful when I do square roots is to write the prime factors and group up pairs. \[\frac{16}{36}=\frac{2*2*2*2}{2*2*3*3}=\]

OpenStudy (anonymous):

let them learn

OpenStudy (anonymous):

That was me, I am trying to learn this

OpenStudy (anonymous):

I do not understand the rule

OpenStudy (anonymous):

ok by dividing both numerator and denominator by 4 you got 4 and 9

OpenStudy (anonymous):

yes I did

OpenStudy (anonymous):

apply the rule on \[\sqrt{\frac{ 4 }{ 9}}=??\]

OpenStudy (anonymous):

carefully look at the rule..

OpenStudy (anonymous):

4 will not equally divide by 9 so that is it simplified

OpenStudy (anonymous):

yes you cant simplify any further so apply the rule

OpenStudy (anonymous):

then it cannot be squared. I am not sure of the rule you are talking about.

OpenStudy (anonymous):

\[\sqrt{\frac{ 16 }{ 36 }}=\sqrt{\frac{ 4 }{ 9 }}=\frac{ \sqrt{4} }{ \sqrt{9} }\]

OpenStudy (anonymous):

can u do the final step now???

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

by the way, the final answer is 2 --- 3

OpenStudy (anonymous):

LOL I already figured tha tout.

OpenStudy (anonymous):

good for you....☺